OpenStudy (anonymous):

solve integral cos^2(2x)

OpenStudy (anonymous):

\[\int\limits \cos^2(2x)\]

OpenStudy (anonymous):

dx

hartnn (hartnn):

you know the relation for cos 2x ?

OpenStudy (anonymous):

i use: 1+cos2x/2

OpenStudy (anonymous):

that is how it's being solved in my textbook

hartnn (hartnn):

ok, so cos^2 2x will be (1+cos 4x)/2 right ?

OpenStudy (anonymous):

then the answer should be \[\frac{ x }{ 2 } + \frac{ \sin2x }{ 4 }\] + k

OpenStudy (anonymous):

oh, it's cos4x/2?

hartnn (hartnn):

2nd term = cos 4x/2 its integral = (1/2) sin 4x /4 ok ?

hartnn (hartnn):

yes, 2nd term is (cos 4x )/2 any doubts ?

OpenStudy (anonymous):

yes, i'm checking

OpenStudy (anonymous):

it appears my professor wrote it wrong

OpenStudy (anonymous):

and i've been stuck for hours

hartnn (hartnn):

2nd term = cos 4x/2 its integral = (1/2) sin 4x /4 got this ?

OpenStudy (anonymous):

yes yes, i got it now

hartnn (hartnn):

good :) ask if any more doubts anywhere ...

OpenStudy (anonymous):

nope, i'll try and solve it again and thank you!

hartnn (hartnn):

welcome ^_^